# Theodore S. Rappaport, Mingjun Ying, Nicola Piovesan, Antonio De Domenico, and Dipankar Shakya, “Waste Factor and Waste Figure: A Unified

### v1

@@ II-C Waste Factor for a cascaded system
$P_{\text{consumed,path1}}$ in eq. (8) includes the input power from the source output and also comprises both the useful signal power transmitted to the subsequent device and the power wasted by device 1 itself. When $P_{\text{source,out}}$ is subtracted, we arrive at the standalone signal-path power consumption of device 1 ($D1$).
@@
The total power consumption on the signal path can be expressed by (II-C), which is the sum of power consumed by each device and the signal power input (e.g., signal source output power) of the system.
$P_{\text{consumed,path}}$ $=P_{\text{consumed,path}}^{D1}+P_{\text{consumed,path}}^{D2}$
$+P_{\text{source,out}}.$
@@
$P_{\text{signal,2}}=G_{2}P_{\text{signal,1}},$ (11)
@@
$W=\left(W_{2}+\frac{W_{1}-1}{G_{2}}\right).$ (13)
@@
Despite the similarity of the math for $F$ and $W$, [2, 1], it is worth noting that $F$ concerns additive noise, which is of primary consequence in the design of a receiver, whereas $W$ concerns the additive wasted power which, as shown in this paper, has broad application in many realms wherever there is a transmission of information from a source to a sink. For example, it is shown in [2] and in Section V that the Waste Factor of an attenuator is simply its loss, and the Waste Factor of any channel is simply the channel loss. These principles are used subsequently.
@@ III Waste Factor Analysis for Parallel Communication Systems
MISO, SIMO, and MIMO are prevalent architectures in wireless networks, known for enhancing system performance, capacity, and reliability. This section introduces the concept of Waste Factor for paralleled systems and outlines a comprehensive methodology for evaluating the power efficiency of MISO, SIMO, and MIMO configurations. By analyzing the efficiencies and gains of individual components within these paralleled chains, we establish a generalized theory to compute the overall $W$ of these systems. The analysis provided herein deepens the understanding of using Waste Factor for analyzing power efficiency in complex multi-antenna or multi-user architectures.
@@ III-A W for Multiple-Input Single-Output (MISO) System
To investigate the MISO structure, we first consider a configuration consisting of two transmitters (TXs) and one receiver (RX). In this setup, each TX communicates with the RX via uncorrelated channels. These channels are characterized by zero-mean additive white Gaussian noise (AWGN).
@@ III-A1 W for non-coherent combining MISO
In Fig. 3, TX1 is characterized by a Waste Factor $W_{\text{T1}}$, operating over a channel with $W_{\text{C1}}=L_{\text{C1}}=1/G_{\text{C1}}$, and delivering received power $P_{1}$ into RX1, which has its own Waste Factor $W_{\text{R1}}$ and receiver gain $G_{\text{R1}}$. Similarly, TX2 has Waste Factor $W_{\text{T2}}$ and transmits to RX1 with channel loss of $L_{\text{C2}}=1/G_{\text{C2}}$ which yields $W_{\text{C2}}$ and received power $P_{2}$ to RX1. According to the results of a lossy channel study in Section V and [2], the Waste Factor $W$ for each of these two parallel systems applied to RX1 can be considered. When assuming non-coherent combining at the input of RX1, as discussed in [34], the received powers $P_{1}$ and $P_{2}$ are combined in the input to the last stage of the MISO system (shown as a dashed box in Fig. 3), where the Waste Factor of the system just before the last stage of the MISO system (e.g., just before the receiver in Fig. 3) is defined by a single Figure of Merit $W_{\text{2}\parallel}$. For two received signals $r_{1}(t)$ and $r_{2}(t)$ with powers $P_{1}$ and $P_{2}$ respectively, non-coherent combining involves no phase information and assumes that the phases of incoming signals from each transmitter are uniformly and identically distributed (i.i.d.). Therefore, the combined receiver power average in non-coherent combining at the RX antenna is simply the sum of individual powers: $P_{\text{signal,noncoh}}=P_{1}+P_{2}$ [34]. Then, using eqs. (3), (12), and (II-C), the Waste Factor for a MISO system is given by
@@
$=\frac{{P_{1}\left(W_{\text{C1}}+\frac{W_{\text{T1}}-1}{G_{\text{C1}}}\right)}+{P_{2}\left(W_{\text{C2}}+\frac{W_{\text{T2}}-1}{G_{\text{C2}}}\right)}}{P_{1}+P_{2}}$ (18)
@@ III-A2 W for coherent combining MISO
$W_{\text{M}\mathbin{\parallel}}^{\text{coh}}=\frac{\sum_{i=1}^{M}\left({P_{N,i}\times W_{\text{cascade},i}}\right)}{\left|\sum_{i=1}^{M}{\sqrt{P_{N,i}}}\right|^{2}}=\frac{\sum_{i=1}^{M}\left({\gamma_{i}W_{\text{cascade},i}}\right)}{\left|\sum_{i=1}^{M}{\sqrt{\gamma_{i}}}\right|^{2}}$ (25)
@@
$W_{\text{MISO}}=W_{N+1}+\frac{(W_{\text{M}\mathbin{\parallel}}-1)}{G_{N+1}}.$ (26)
@@ III-C1 W for non-coherent combining MIMO
$\begin{bmatrix}P_{\text{R1}}^{\text{noncoh}}&P_{\text{R2}}^{\text{noncoh}}\end{bmatrix}=\begin{bmatrix}P_{\text{T1}}&P_{\text{T2}}\end{bmatrix}\begin{bmatrix}W_{\text{C11}}^{-1}&W_{\text{C12}}^{-1}\\ W_{\text{C21}}^{-1}&W_{\text{C22}}^{-1}\end{bmatrix}.$ (27)
@@ III-C3 W for general MIMO system
$W_{\text{MINO}}=W_{\text{N}\mathbin{\parallel}}+\frac{(W_{\text{MINO}}^{1}-1)}{G_{\text{N}\mathbin{\parallel}}}.$ (39)
@@ III-D Examples of Waste Factor in practical systems
Some applications of Waste Factor are given in [2], where the calculations and use of $W$ are demonstrated for making energy-efficient decisions and for computing actual power consumption levels in circuits, systems, channels, and data centers. Section V expands the practical use of the Waste Factor theory to the design of RAN components.
@@ IV Radio Access Network (RAN) Analysis Using Waste Factor
IV-A Need for unified energy efficiency metric
The advent of 5G wireless technology marks a significant leap forward, introducing innovative services, technologies, and networking paradigms. These advancements offer substantial societal benefits, ranging from enhanced mobile broadband to private networks with mission-critical communications. Despite these advancements, the deployment of 5G networks raises concerns regarding energy consumption. While 5G networks have been found to be approximately four times more energy-efficient per unit of data than their 4G counterparts, their overall energy consumption remains substantially higher[35]. This increase is attributed to the need for more power-hungry processing at each site to accommodate wider bandwidths and greater numbers of radio chains per base station, as well as the need for more cell sites to achieve comparable coverage at higher frequencies, i.e., C-band, mid-band, and millimeter wave bands [10, 32, 36, 31].
The financial implications of increased energy consumption are significant, particularly for mobile network operators. Network costs, which on average constitute about 25% of the operational expenditure (OPEX) of a mobile network operator, are heavily impacted by energy expenses, with approximately 90% of network costs dedicated to energy bills[37]. A nationwide carrier with 100 million subscribers pays about \$2 Billion/year for electricity, and when multiplied over many hundreds of global carriers and 4 billion subscribers, the electricity bill for the cellular industry is approximately \$1 Trillion/year.
A critical examination reveals that the majority of energy consumption—over 76%—occurs within the radio access network (RAN), as shown in Section I, highlighting the need for energy-efficient strategies focused on the RAN components of the network.
The challenge of reducing network energy consumption is compounded by the anticipated surge in mobile data traffic, expected to increase by a factor of 3 between 2023 and 2029 [38]. Meeting this demand will necessitate the deployment of additional hardware with wider channel bandwidths, further exacerbating energy consumption issues. The trajectory of this growth underscores the urgent need for sustainable network expansion strategies and accurate, standardized energy efficiency metrics.
In response to the issues of increased energy consumption and the need for sustainable networks, the 3GPP has initiated efforts to mitigate RAN energy consumption. The 3GPP new radio (NR) Release 15 has introduced intra-NR network energy-saving solutions, including autonomous cell switch-off and re-activation features for capacity-boosting cells through Xn/X2 interfaces. Furthermore, 3GPP NR Release 17 has specified inter-system network energy-saving solutions and recognized network energy saving as a potential application for artificial intelligence. Despite efforts in addressing the energy consumption issue in the network, the effectiveness of current 3GPP long term evolution (LTE) and 5G NR energy-saving solutions fall short of addressing the fundamental energy consumption challenges [39].
Given the rising energy demands of the RAN, it becomes imperative to develop and implement a robust metric for measuring energy efficiency. Such a metric would facilitate the evaluation of energy usage in relation to network performance, enabling more informed decisions aimed at reducing the energy footprint of RAN operations. In response to the need for a new sustainability metric, this section applies W as a novel and effective approach for analyzing energy utilization within the RAN, and shows why a new metric such as W is needed urgently.
IV-B RAN Architecture
@@
IV-C RAN Energy Efficiency Metrics
@@
IV-D Problems with the Standard RAN EE Metrics
@@
Despite being less efficient in transmitting data, BS-B is evaluated as more energy-efficient (e.g., having a higher data volume per energy) than BS-A according to the EE metric in eq. (42). This EE gives a wrong result, since if BS-A was more heavily loaded, it would be much more energy efficient. It is clear that (42) is dependent on the particular loading and not the power efficiency of the equipment. This problem occurs because BS-A handles low data volume and, in this condition, the total consumed power is largely impacted by the non-path power consumption, which is not related to the actual energy efficiency of source-to-sink transmission. Note that in such a case, the actual transmission efficiency is obscured due to the particular power level (e.g. loading) within the definition of EE. Conversely, BS-B handles a large data volume, which dilutes the impact of its large per-GB data transmission energy cost.
@@
Finally, W is proposed as a unified metric to assess the end-to-end transmission and reception from the energy perspective, including the impact of the wireless channel and of the user equipment, thereby allowing a unified, harmonious energy efficiency metric for the network, site, and RU components over widely varying operating conditions and propagation environments.
@@
IV-E Waste Factor of Passive Components
@@
$=\frac{P_{\text{in}}-P_{\text{out}}+P_{\text{out}}}{P_{\text{out}}}=\frac{P_{\text{in}}}{P_{\text{out}}}=L,$ (47)
Eq. (IV-E) underscores how $W$ is equivalent to the loss factor $L$ of the passive element. In essence, W directly reflects the efficiency loss in passive elements, indicating that for passive components, W and the loss are identical concepts [2].
IV-E1 Passive RF Mixers
@@ IV-E3 Antenna
$\Gamma=\frac{Z_{\text{Ant}}-Z_{\text{out}}}{Z_{\text{Ant}}+Z_{\text{out}}}=\frac{\mathrm{VSWR}-1}{\mathrm{VSWR}+1},$ (52)
@@
IV-F Waste Factor of Active Components
@@ IV-F3 Power Amplifier
$\text{PAE}^{\#2}=\frac{P_{\text{out}}-P_{\text{in}}}{P_{\text{DC}}},$ (63)
@@ IV-F4 Low Noise Amplifier
$=\frac{G}{\text{FoM}_{\text{LNA}}\text{SNR}_{\text{in}}(F-1)GN_{\text{in}}}$ (66)
$=\frac{G}{\text{FoM}_{\text{LNA}}\text{SNR}_{\text{in}}P_{\text{additive-noise}}}.$ (67)
where $\text{SNR}_{\text{in}}=P_{\text{in}}/N_{\text{in}}=P_{\text{out}}/(GN_{\text{in}})$ denotes the input Signal-to-Noise Ratio. Typically, $W_{\text{LNA}}$ approaches unity under standard operational conditions, as shown in eq. (IV-F4). It is important to note that the power consumption of LNAs is generally minuscule when compared to other components, and the large gain of LNAs, can be shown in eq. (II-C) to minimize the impact of overall W of cascade components when moving closer to the source [2].
IV-G Analyzing Quiescent Power in Amplifiers

### v2

@@ II-C Waste Factor for a cascaded system
where $P_{\text{consumed,path1}}$ in eq. (8) includes the input power from the source output and also comprises both the useful signal power transmitted to the subsequent device and the power wasted by device 1 itself. When $P_{\text{source,out}}$ is subtracted, we arrive at the standalone signal-path power consumption of device 1 ($D1$).
@@
The total power consumption on the signal path can be expressed by (11), which is the sum of power consumed by each device and the signal power input (e.g., signal source output power) of the system.
$P_{\text{consumed,path}}=P_{\text{consumed,path}}^{D1}+P_{\text{consumed,path}}^{D2}+P_{\text{source,out}}.$ (11)
@@
$P_{\text{signal,2}}=G_{2}P_{\text{signal,1}}.$ (12)
@@
$W=\left(W_{2}+\frac{W_{1}-1}{G_{2}}\right),$ (14)
@@
Despite the similarity of the math between $F$ and $W$, [28, 6], it is worth noting that $F$ concerns additive noise, which is of primary consequence in the design of a receiver, whereas $W$ concerns the additive wasted power which, as shown in this paper, has broad application in many realms wherever there is a transmission of information from a source to a sink. For example, it is shown in [28] and in Section IV that the Waste Factor of an attenuator is simply its loss, and the Waste Factor of any channel is simply the channel loss. These principles are used subsequently.
@@ III Waste Factor Analysis for Parallel Communication Systems
MISO, SIMO, and MIMO are prevalent architectures in wireless networks, known for enhancing system performance, capacity, and reliability. This section introduces the concept of Waste Factor for paralleled systems and outlines a comprehensive methodology for evaluating the power efficiency of MISO, SIMO, and MIMO configurations. By analyzing the efficiencies and gains of individual components within these paralleled chains, we establish a generalized theory to compute the overall $W$ of these systems. The analysis provided herein expands the utility of Waste Factor for analyzing power efficiency in complex multi-antenna and multi-user system and network architectures.
@@ III-A W for Multiple-Input Single-Output (MISO) System
To investigate the MISO structure, we first consider a configuration consisting of two transmitters (TXs) and one receiver (RX). In this setup, each TX communicates with the RX via uncorrelated channels. These channels are assumed to be zero-mean additive white Gaussian noise (AWGN).
@@ III-A1 W for non-coherent combining MISO
In Fig. 3, TX1 is characterized by a Waste Factor $W_{\text{T1}}$, operating over a channel with Waste Factor $W_{\text{C1}}=L_{\text{C1}}=1/G_{\text{C1}}$, (see [28] and Section IV-G)) and delivering received power $P_{1}$ into RX1, which has its own Waste Factor $W_{\text{R1}}$ and receiver gain $G_{\text{R1}}$. Similarly, TX2 has Waste Factor $W_{\text{T2}}$ and transmits to RX1 with channel loss of $L_{\text{C2}}=1/G_{\text{C2}}$ which yields $W_{\text{C2}}$ and received power $P_{2}$ to RX1. According to the results of a lossy channel study in Section IV and [28], the Waste Factor $W$ for each of these two parallel systems applied to RX1 can be considered. When assuming non-coherent combining at the input of RX1, as discussed in [38], the received powers $P_{1}$ and $P_{2}$ are combined in the input to the last stage of the MISO system (shown as a dashed box in Fig. 3), where the Waste Factor of the system just before the last stage of the MISO system (e.g., just before the receiver in Fig. 3) is defined by a single Figure of Merit $W_{\text{2}\parallel}$. For two received signals $r_{1}(t)$ and $r_{2}(t)$ with powers $P_{1}$ and $P_{2}$ respectively, non-coherent combining involves no phase information and assumes that the phases of incoming signals from each transmitter are uniformly and identically distributed (i.i.d.). Therefore, the combined receiver power average in non-coherent combining at the RX antenna is simply the sum of individual powers: $P_{\text{signal,noncoh}}=P_{1}+P_{2}$ [38]. Then, using eqs. (3), (13), and (II-C), the Waste Factor for a MISO system is given by
@@
$=\frac{{P_{1}\left(W_{\text{C1}}+\frac{W_{\text{T1}}-1}{G_{\text{C1}}}\right)}+{P_{2}\left(W_{\text{C2}}+\frac{W_{\text{T2}}-1}{G_{\text{C2}}}\right)}}{P_{1}+P_{2}}.$ (19)
@@ III-A2 W for coherent combining MISO
$W_{\text{M}\mathbin{\parallel}}^{\text{coh}}=\frac{\sum_{i=1}^{M}\left({P_{N,i}\times W_{\text{cascade},i}}\right)}{\left|\sum_{i=1}^{M}{\sqrt{P_{N,i}}}\right|^{2}}=\frac{\sum_{i=1}^{M}\left({\gamma_{i}W_{\text{cascade},i}}\right)}{\left|\sum_{i=1}^{M}{\sqrt{\gamma_{i}}}\right|^{2}}.$ (26)
@@
$W_{\text{MISO}}=W_{N+1}+\frac{(W_{\text{M}\mathbin{\parallel}}-1)}{G_{N+1}},$ (27)
@@ III-B W for Single-Input Multiple-Output (SIMO) System
Additionally, this analysis exploits the fundamental Waste Factor theory given in Section  II, and quantifies the cumulative impact of Waste Factors when cascades are combined using MISO or SIMO structures. W for any device remains constant if each component operates within its linear range, regardless of the transceiver implementation. For instance, in OFDM systems, linearity is ensured through backoff to prevent distortion. The linearity of each component is also required for the Waste Factor theory to be applied to single carrier and carrier aggregation scenarios, which involve specific hardware components in the cascade to determine overall system energy efficiency.
@@ III-C1 W for non-coherent combining MIMO
$\begin{bmatrix}P_{\text{R1}}^{\text{noncoh}}&P_{\text{R2}}^{\text{noncoh}}\end{bmatrix}=\begin{bmatrix}P_{\text{T1}}&P_{\text{T2}}\end{bmatrix}\begin{bmatrix}W_{\text{C11}}^{-1}&W_{\text{C12}}^{-1}\\ W_{\text{C21}}^{-1}&W_{\text{C22}}^{-1}\end{bmatrix},$ (28)
@@ III-C3 W for general MIMO system
$W_{\text{MINO}}=W_{\text{N}\mathbin{\parallel}}+\frac{(W_{\text{MINO}}^{1}-1)}{G_{\text{N}\mathbin{\parallel}}},$ (40)
@@ III-D Examples of Waste Factor in practical systems
Several practical applications of Waste Factor are detailed in [28], where calculations and usage of $W$ are demonstrated for making energy-efficient decisions and to accurately determine power consumption in circuits, systems, channels, and data centers. The exploration of $W$ in MIMO structures notably broadens its applicability, particularly in advanced systems like emerging cell-free massive MIMO and ultra-dense networks. When the input and output structures of these networks are well-defined, applying the Waste Factor is straightforward by referring to the method proposed in Section III-A, B, and C. This method can be applied in general to any wireless or wired communication systems with parallel structures, as discussed in Section IV.
In the following sections, we apply the Waste Factor theory for cascaded systems (presented in Section II) and the Waste Factor theory for parallel systems (derived in Section III-A, B, and C) to the RAN scenario. In particular, Section IV introduces the RAN architecture and derives the Waste Factor for the main cascaded components in any RAN. Section V provides analysis and simulation results of a RAN consisting of distributed MU-MIMO networks, showcasing the benefits of applying the Waste Factor theory to assess and enhance the energy efficiency of RANs through real-world system-based power and energy consumption models.
@@ IV Radio Access Network (RAN) Analysis Using Waste Factor
Given the rising energy demands of the RAN, it becomes imperative to develop and implement a robust metric for measuring energy efficiency. In response to the need for a new sustainability metric, this section applies W as a novel and effective approach for analyzing energy utilization within the RAN, and shows why a new metric such as W is needed urgently.
IV-A RAN Architecture
@@
As described in [39], a typical RAN site includes a variety of equipment essential for mobile communications. It is important to recognize that RAN architecture can vary significantly, ranging from traditional setups to more modern configurations like Cloud RAN (C-RAN), Distributed RAN (D-RAN), Virtualized RAN (vRAN), and Open RAN (O-RAN). Each of these architectures offers unique benefits and challenges in terms of deployment and energy consumption.
A key observation is that all RAN architectures, and in fact, virtually all practical communication systems, require linear signal replication over huge orders of magnitude of power levels between a source to sink so that signal fidelity and integrity are preserved and so that a receiver may differentiate between digital waveforms. The Waste Factor, W, provides a universal metric capable of assessing total power consumption across all forms of RAN architectures. By applying W, network operators can evaluate and compare the energy efficiency of different RAN setups under a standardized framework, enabling targeted improvements and optimizations. Moreover, W is particularly valuable in identifying and quantifying the most energy-hungry components or processes within each RAN structure, guiding the development of more sustainable network solutions.
IV-B RAN Energy Efficiency Metrics
@@
In this section, we overview the previous work carried out by global standard bodies in defining energy efficiency metrics for the RAN both as a cascaded system and for individual RAN components along the cascade.
@@
Importantly, as we demonstrate in Section IV and V, all of these widely adopted standardized energy efficiency metrics suffer a major issue that prevents them from being effectively used to assess and compare the energy efficiency of the RAN and the impact of components in transmitting or receiving data.
IV-C Problems with the Standard RAN EE Metrics
@@
Despite being less efficient in transmitting data, BS-B is evaluated as more energy-efficient (e.g., having a higher data volume per energy) than BS-A according to the EE metric in eq. (43).
This EE metric produces a misleading result because it depends on the traffic load at which it is measured rather than on the power efficiency of the equipment. For example, if BS-A were more heavily loaded, it would appear to be much more energy efficient. In more detail, the issue arises because BS-A handles a low data volume, and under such conditions, the total consumed power is significantly influenced by the non-path power consumption, which is not related to the actual energy efficiency of source-to-sink transmission. Consequently, the true transmission efficiency is obscured by the specific power level (e.g., loading) considered in the EE metric. In contrast, BS-B handles a large data volume, which dilutes the relative impact of its substantial per-GB data transmission energy cost.
@@
Finally, W is proposed as a unified metric to assess the end-to-end transmission and reception from the energy perspective, including the impact of the wireless channel and of the user equipment, thereby allowing a unified, harmonious energy efficiency metric for the network, site, and RU components over widely varying operating conditions and propagation environments. To this end, in the following subsections, we use the fundamentals of the Waste Factor theory to derive the Waste Factor for all passive and active components of the RAN, followed by the derivation of the Waste Factor for RUs and UEs.
@@
IV-D Waste Factor of Passive Components
@@
$=\frac{P_{\text{in}}-P_{\text{out}}+P_{\text{out}}}{P_{\text{out}}}=\frac{P_{\text{in}}}{P_{\text{out}}}=L.$ (48)
Eq. (IV-D) underscores how $W$ is equivalent to the loss factor $L$ of the passive element. In essence, W directly reflects the efficiency loss in passive elements, indicating that for passive components, W and the loss are identical concepts [28].
IV-D1 Passive RF Mixers
@@ IV-D3 Antenna
$\Gamma=\frac{Z_{\text{Ant}}-Z_{\text{out}}}{Z_{\text{Ant}}+Z_{\text{out}}}=\frac{\mathrm{VSWR}-1}{\mathrm{VSWR}+1}.$ (53)
@@
IV-E Waste Factor of Active Components
@@ IV-E3 Power Amplifier
$\text{PAE}^{\#2}=\frac{P_{\text{out}}-P_{\text{in}}}{P_{\text{DC}}}.$ (64)
@@ IV-E4 Low Noise Amplifier
$=\frac{G}{\text{FoM}_{\text{LNA}}\text{SNR}_{\text{in}}(F-1)GN_{\text{in}}}$
$=\frac{G}{\text{FoM}_{\text{LNA}}\text{SNR}_{\text{in}}P_{\text{additive-noise}}},$ (67)
where $\text{SNR}_{\text{in}}=P_{\text{in}}/N_{\text{in}}=P_{\text{out}}/(GN_{\text{in}})$ denotes the input Signal-to-Noise Ratio. Typically, $W_{\text{LNA}}$ approaches unity under standard operational conditions, as shown in eq. (IV-E4). It is important to note that the power consumption of LNAs is generally minuscule when compared to other components, and the large gain of LNAs, can be shown in eq. (II-C) to minimize the impact of overall W of cascade components when moving closer to the source [28].
IV-F Analyzing Quiescent Power in Amplifiers